# Decoupled Split Learning via Auxiliary Loss

### v1

@@ I Introduction
More specifically, a canonical SL setup involves splitting a neural network (in total $L$ layers) into two parts: a client-side model (from layer $1$ to $L_{c}$) on the client side and a server-side model (from layer $L_{c}+1$ to $L$) on the server side. Layer $L_{c}$ in which the network is split is called the cut layer (see also Fig. 1). During a forward pass, the client computes intermediate activations and sends them (and the labels) to the server. [^1] Sharing labels but not raw features is widely considered in SL. One can adopt a U-shaped SL structure that does not share labels either [4]. The server continues the forward propagation to compute predictions. In the backward pass, the server evaluates the loss and sends the gradient of the intermediate activation back to the client, which then updates the clients-side model’s parameters using the chain rule. This approach keeps data local and only exchanges intermediate features and gradients, which helps preserve privacy. It also enables less heavy client-side computation (compared to other distributed learning approaches such as federated learning), since the clients only need to train part of the model.
@@ IV-B Experiment Results
Convergence and Performance (Q1). Figure 4 plots the test accuracy over training epochs. We observe that CSL and DSL exhibit similar convergence behavior on CIFAR-10 and CIFAR-100, both achieving accuracies that are comparable across different numbers of clients $N$ and model split strategies $L_{c}$. The learning curves indicate that our decoupling approach does not necessarily introduce optimization instability, and the performance gap between CSL and DSL remains small during training. In addition, across different model splits $L_{c}$, we observe that deeper cuts tend to achieve better performance (e.g., DSL with $N=5$), which is consistent with [3].
@@
Communication (Q2). Figure 5 reports the communication volume breakdown between forward and backward transmissions under different model splits $L_{c}$ and N=10N=10.333Note that results are invariant to client number $N$, as the communication volume depends on the size of the cut layer. When the cut layer is placed early in ResNet-110, CSL incurs substantial backward communication due to the transmission of gradients, which results in a total communication volume of approximately 6.5 GB over training on both datasets. In contrast, DSL eliminates backward communication by design, and it reduces the communication volume by nearly 50% while preserving the same forward communication cost. As the cut layer moves toward the deep, the overall communication volume decreases for both methods due to the reduced cut layer dimensionality.
@@
Memory Usage (Q2). Figure 6 reports the peak memory usage under different client numbers $N$ and model splits $L_{c}$. For CSL, the peak memory footprint remains largely constant across different choices of $L_{c}$, as BP requires storing intermediate activations regardless of where the cut layer is placed. In contrast, DSL exhibits a clear reduction in peak memory usage as the cut layer moves shallower into the network. When $L_{c}=s$, DSL reduces the peak memory consumption by up to $58\%$ compared to CSL on both CIFAR-10 and CIFAR-100. This reduction stems from the elimination of activation storage needed for BP, with earlier cuts further limiting the total size of activations that must be retained at the client. The observed trends are consistent across datasets and numbers of clients, which shows that DSL provides substantial and robust memory savings without sacrificing model performance.

### v2

@@ I Introduction
More specifically, a canonical SL setup involves splitting a neural network (in total $L$ layers) into two parts: a client-side model (from layer 11 to $L_{c}$) on the client side and a server-side model (from layer $L_{c}+1$ to $L$) on the server side. Layer $L_{c}$ in which the network is split is called the cut layer (see also Fig. 1). During a forward pass, the client computes intermediate activations and sends them (and the labels) to the server. [^1] Sharing labels but not raw features is widely considered in SL. One can adopt a U-shaped SL structure that does not share labels either [4]. The server continues the forward propagation to compute predictions. In the backward pass, the server evaluates the loss and sends the gradient of the intermediate activation back to the client, which then updates the clients-side model’s parameters using the chain rule. This approach keeps data local and only exchanges intermediate features and gradients, which helps preserve privacy. It also enables less heavy client-side computation (compared to other distributed learning approaches such as federated learning), since the clients only need to train part of the model.
@@ IV-B Experiment Results
Convergence and Performance (Q1). Figure 4 plots the test accuracy over training epochs. We observe that CSL and DSL exhibit similar convergence behavior on CIFAR-10 and CIFAR-100, both achieving accuracies that are comparable across different numbers of clients NN and model split strategies $L_{c}$. The learning curves indicate that our decoupling approach does not necessarily introduce optimization instability, and the performance gap between CSL and DSL remains small during training. In addition, across different model splits $L_{c}$, we observe that deeper cuts tend to achieve better performance (e.g., DSL with $N=5$), which is consistent with [3].
@@
Communication (Q2). Figure 5 reports the communication volume breakdown between forward and backward transmissions under different model splits $L_{c}$ and N=10N=10.333Note that results are invariant to client number NN, as the communication volume depends on the size of the cut layer. When the cut layer is placed early in ResNet-110, CSL incurs substantial backward communication due to the transmission of gradients, which results in a total communication volume of approximately 6.5 GB over training on both datasets. In contrast, DSL eliminates backward communication by design, and it reduces the communication volume by nearly 50% while preserving the same forward communication cost. As the cut layer moves toward the deep, the overall communication volume decreases for both methods due to the reduced cut layer dimensionality.
@@
Memory Usage (Q2). Figure 6 reports the peak memory usage under different client numbers NN and model splits $L_{c}$. For CSL, the peak memory footprint remains largely constant across different choices of $L_{c}$, as BP requires storing intermediate activations regardless of where the cut layer is placed. In contrast, DSL exhibits a clear reduction in peak memory usage as the cut layer moves shallower into the network. When $L_{c}=s$, DSL reduces the peak memory consumption by up to $58\%$ compared to CSL on both CIFAR-10 and CIFAR-100. This reduction stems from the elimination of activation storage needed for BP, with earlier cuts further limiting the total size of activations that must be retained at the client. The observed trends are consistent across datasets and numbers of clients, which shows that DSL provides substantial and robust memory savings without sacrificing model performance.